equation of hyperbola is given by the equation 8y2-2x2=16. find the co ordinates of.
1) foci
2) vertices
3) eccentricity

Answer by lwsshak3(11628)

equation of hyperbola is given by the equation 8y2-2x2=16. find the co ordinates of.
1) foci
2) vertices
3) eccentricity
***
8y^2-2x^2=16
y^2/2-x^2/8=1
This is an equation of a hyperbola with vertical transverse axis and center at (0,0)
Its standard form of equation: y^2/a^2-x^2/b^2=1
..
a^2=2
a=√2
b^2=8
b=√8
c^2=a^2+b^2
c^2=2+8=10
c=√10≈3.16
..
1) foci:(0,0±c)=(0,0±3.16)=(0,0-3.16) and (0,0+3.16)
2) vertices:(0,0±a)=(0,0±√2)=(0,0-√2)and (0,0+√2)
3) eccentricity: c/a=√10/√2=√5